OpenStudy (anonymous):

If someone has driven 24 miles and that is two thirds of the drive then what is the total length

OpenStudy (kt7779):

take half of that plus the original!

OpenStudy (kt7779):

so 12 + 24= 36

OpenStudy (anonymous):

2/3 x = 24 Solve for x

OpenStudy (kt7779):

1/2 of 2/3 is 1/3 and 1/3 + 2/3 = 3/3 = 1

OpenStudy (anonymous):

(2/3)/24 = (1/3)/x cross multiply 2/3x = (24)(1/3) 2/3x = 8 x = 8 divided by 2/3 x = 8 x 3/2 x = 24/2 x = 12 I just figured that 1/3 = 12 so 2/3 = 24.....making 3/3 = 36 total length is 36 miles

OpenStudy (anonymous):

2/3 x = 24 By dividing both sides by 2/3 you get x=36

OpenStudy (anonymous):

that will work as well :)

OpenStudy (anonymous):

I think it's the whole point to algebra, isnt it? :-)

OpenStudy (anonymous):

yes it is....so is mine

OpenStudy (anonymous):

Thank everyone but I have the answer now kt7779 gave me the answer I needed.!!!

OpenStudy (kt7779):

your welcome!

OpenStudy (anonymous):

but you got the right answer...thats the point

OpenStudy (anonymous):

I guess so.. Thanks to everyone who gave me the answer!!!!

OpenStudy (anonymous):

thats what we are here for.....not to give the answer, but to help...although we did just give you the answer...lol

OpenStudy (kt7779):

well im glad we could all help her and show her in different ways so that she can choose what helps her the most and she can decide whatever one she understands most. that's shy im here!

OpenStudy (kt7779):

why